OpenStudy (123ab456c):

I BET NO ONE CAN ANSWER THIS MATH PROBLEM E= LET SEE HOW MANY PEOPLE WILL TRY TO FIGURE IT OUT EVERY ONE COME TO MY QUESTION!!!

OpenStudy (studygurl14):

2.71828182845904523536028747135266249775724709369995...

OpenStudy (solomonzelman):

More precise, \(\color{#000000 }{ \displaystyle e^x=\sum_{n=0}^{\infty}\frac{x^n}{n!} \quad \Rightarrow \quad e^1=\sum_{n=0}^{\infty}\frac{1^n}{n!} \quad \Rightarrow \quad e=\sum_{n=0}^{\infty}\frac{1}{n!}}\)

OpenStudy (solomonzelman):

but, there is a lot of that stuff on wiki, or other online sources all over the place.

OpenStudy (123ab456c):

you are all WRONG! keep tryin:):) youll get it